# A Quantum-Driven Evolutionary Framework for Solving High-Dimensional Sharpe Ratio Portfolio Optimization

### v1

@@ Abstract.
High-dimensional portfolio optimization faces significant computational challenges under complex constraints, with traditional optimization methods struggling to balance convergence speed and global exploration capability. To address this, firstly, we introduce an enhanced Sharpe ratio–based model that incorporates all constraints into the objective function using adaptive penalty terms, transforming the original constrained problem into an unconstrained single-objective formulation. This approach preserves financial interpretability while simplifying algorithmic implementation. To efficiently solve the resulting high-dimensional optimization problem, we propose a Quantum Hybrid Differential Evolution (QHDE) algorithm, which integrates Quantum-inspired probabilistic behavior into the standard DE framework. QHDE employs a Schrödinger-inspired probabilistic mechanism for population evolution, enabling more flexible and diversified solution updates. To further enhance performance, a good point set–chaos reverse learning strategy is adopted to generate a well-dispersed initial population, and a dynamic elite pool combined with Cauchy-Gaussian hybrid perturbations strengthens global exploration and mitigates premature convergence. Experimental validation on CEC benchmarks and real-world portfolios involving 20 to 80 assets demonstrates that QHDE’s performance improves by up to 73.4%. It attains faster convergence, higher solution precision, and greater robustness than seven state-of-the-art counterparts, thereby confirming its suitability for complex, high-dimensional portfolio optimization and advancing quantum-inspired evolutionary research in computational finance.
Portfolio optimization, Differential Evolution, Sharpe ratio-based portfolio, Quantum evolution strategy, Dynamic elite pool
copyright: acmlicensedjournalyear: 2026conference: ; $c$s: Computing methodologies Discrete space search
1. Introduction
@@
In finance, a portfolio represents a collection of stocks or assets managed by individuals or institutions. Portfolio optimization entails distributing a fixed budget among the assets to minimize risk while maximizing expected returns. The Modern Portfolio Theory (MPT) employs the mean-variance model to describe the relationship between risk and return (Markowitz, 1952). The core principle of this theory quantifies risk using variance, where lower-risk portfolios are typically characterized by smaller variances. However, MPT overlooks practical market constraints and the influence of background risks, such as base requirements, upper and lower bounds, and challenges inherent in managing large-scale asset portfolios (Khan et al., 2021; Menoncin, 2002; Tsanakas, 2008). Therefore, real-world portfolio optimization must incorporate these constraints and uncertainties to account for actual market conditions.
@@
Differential Evolution (DE) (Das and Suganthan, 2010) is a widely used metaheuristic algorithm for global optimization, renowned for its high performance and ease of implementation in various optimization tasks. DE has the advantages of being simple to implement, requiring fewer parameters, and possessing strong global search capabilities, widely adopted in various engineering optimization and scientific research tasks (Song et al., 2021; Kumari and Vaisakh, 2022; Song et al., 2023; Zhong et al., 2025; Zhang et al., 2026). Given the hybrid decision-making nature of discrete asset selection and continuous weight allocation in portfolio optimization, the real-valued encoding mechanism of DE facilitates a precise characterization of the continuous search space of asset weights, and combined with specialized mapping strategies to effectively handle complex constraints and multi-objective optimization, thereby adapting to both nonlinear and multi-modal requirements in financial decision-making. However, when confronted with high-dimensional complex portfolio problems, DE faces notable limitations. First, it is prone to becoming trapped in local optima, hindering the search for the global optimum. Second, as the dimensionality increases, the search efficiency declines and convergence slows. These limitations restrict the application effectiveness of DE in the field of portfolio optimization. In response to these challenges, the introduction of quantum computing offers a new breakthrough for improving metaheuristic algorithms.
In recent years, quantum computing has emerged as a transformative computational technology, demonstrating significant potential (Deng et al., 2021; Chen et al., 2020). Quantum computing can exponentially accelerate certain computational tasks through properties such as superposition and entanglement of quantum bits (Luo, 2008; Liu et al., 2016). This acceleration capability confers a distinct advantage in addressing high-dimensional optimization problems with large solution spaces, especially in financial applications such as portfolio optimization, risk assessment, and derivative pricing (Lu and Yang, 2024). In this context, quantum algorithms provide a new perspective on optimization problems, and numerous metaheuristic algorithms based on quantum computing have emerged (Zhu et al., 2024; Brandhofer et al., 2022; Shunza et al., 2023; Singh et al., 2025; Cheng and Sholeh, 2025). Integrating quantum computing with existing optimization approaches can fully harness its global search potential, thereby mitigating the limitations of traditional algorithms under high-dimensional and complex constraints.
@@
In this paper, by combining the strengths and limitations of DE in portfolio optimization with the characteristics of quantum computing, we propose a variant of the DE based quantum theory, referred to as the Quantum Hybrid Differential Evolution Algorithm (QHDE). QHDE incorporates multiple innovative improvements relative to conventional DE. Guided by quantum theory for global search, it more effectively explores optimal portfolio configurations in high-dimensional spaces and converges faster to the global optimum. Furthermore, the incorporation of chaotic reverse learning extends the coverage of initial solutions, resulting in a more advantageous search starting point. Meanwhile, the dynamic elite pool and Cauchy-Gaussian mixed disturbance strategy effectively prevent premature convergence and increase the diversity of the solutions. The combination of these improvement strategies enables QHDE to surpass traditional DE and other metaheuristic algorithms in portfolio optimization, allowing for faster identification of the optimal solution that balances risk and return. The main contributions of this paper are as follows:
(1) Incorporating quantum behavior and probability collapse mechanisms, we design a quantum evolution strategy based on population information exchange, which significantly enhances the algorithm’s global search capability and convergence speed.
@@
(2) We introduce an innovative initialization mechanism that combines the good point set method with chaotic reverse learning. This mechanism enhances the quality of initial solutions and improves algorithmic stability, offering a more efficient starting point for subsequent optimization.
@@
(3) We propose a dynamic elite pool strategy and applies Cauchy– Gaussian random perturbation. This strategy mitigates premature convergence to local optima, enhances solution diversity, and strengthens global search capability.
@@
(4) We apply QHDE to portfolio optimization, where it outperforms seven state-of-the-art algorithms, exceeding the runner-up by up to 73.4%. This approach significantly accelerates convergence while enhancing accuracy, proving to be an efficient and robust solution for high-dimensional problems.
The structure of this paper is as follows: In Section 2, we discuss various modeling approaches for portfolio selection problem and analyze their advantages and disadvantages. In Section 3, we provide a detailed discussion of the original DE algorithm and the improvements proposed in this paper. In Section 4, we evaluate and analyze the performance of QHDE through relevant experiments, and apply the algorithm to solve a portfolio selection problem involving multiple stocks. Finally, we conclude the paper and discusses future research directions in Section 5.
2. Math Equations
@@
This section primarily introduces modeling methods of the portfolio selection problem, laying the foundation for the subsequent introduction of optimization techniques.
2.1. Mean-Variance Model
The Mean-Variance Model serves as a foundational mathematical framework for portfolio optimization. In this model, the objective function focuses on minimizing the portfolio’s overall risk, while the expected mean return is incorporated as a constraint. This ensures the portfolio achieves the desired return level while maintaining minimal risk. Its mathematical form is expressed as follows:
(1) $\min\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}$
(2) $\text{s.t.}\sum_{i=1}^{M}E_{i}\alpha_{i}=R$
(3) $\sum_{i=1}^{M}E_{i}=1$
(4) $0\leq E_{i}\leq 1,\quad i=1,2,\ldots,M$
@@ 2.1. Mean-Variance Model
Equation (1) defines the objective function for minimizing portfolio risk. Here, $M$ denotes the total number of stocks, $E_{i}$ and $E_{j}$ represents the investment weight of the $i$th and $j$th stock respectively, $Q_{ij}$ signifies the covariance between the $i$th and $j$th stocks, and $\alpha_{i}$ denotes the expected return of the $i$th stock. This function aims to minimize the portfolio’s overall risk by optimizing the allocation of weights across the assets. Equation (2) ensures that the total return of the investment portfolio reaches the predetermined expected return $R$. Equation (3) stipulates that the sum of the investment weights for all stocks must equal 1 to guarantee the complete allocation of funds. Equation (4) further limits each investment weight between 0 and 1, which ensures that no short-selling is allowed.
2.2. Efficient Frontier
@@
In portfolio selection problems, investors are required to construct a portfolio that maximizes expected returns while minimizing risk. Unlike the mean-variance model, which treats expected returns as a constraint, the efficient frontier model directly incorporates expected returns into the objective function as one of the key components of optimization. Specifically, the objective function balances the trade-off between risk and return by introducing a penalty term for the return, and the expression is:
(5) $\min\left[\omega\left(\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}\right)-(1-\omega)\left(\sum_{i=1}^{M}E_{i}\alpha_{i}\right)\right]$
(6) $\text{s.t.}\quad\sum_{i=1}^{M}E_{i}=1$
(7) $0\leq E_{i}\leq 1,\quad i=1,2,\ldots,M$
where $\omega\in[0,1]$, and the two constraints in Equation (6) and (7) are consistent with the mean-variance model, ensuring full allocation of funds and non-negativity constraints.
2.3. Sharpe Ratio Model
In the Sharpe Ratio Model, the focus shifts from analyzing these two parameters separately to considering their ratio. The Sharpe ratio measures the excess return per unit of risk and serves as an important indicator that comprehensively reflects the balance between portfolio return and risk. Its mathematical expression is as follows:
(8) $Sharpe-ratio=\frac{R_{p}-R}{std(p)}$
@@ 2.3. Sharpe Ratio Model
(9) $\max\frac{\sum_{i=1}^{M}E_{i}\alpha_{i}-R}{\sqrt{\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}}}$
(10) $\text{s.t.}\quad\sum_{i=1}^{M}E_{i}=1$
(11) $0\leq E_{i}\leq 1,\quad i=1,2,\ldots,M$
2.4. Our unconstrained model
In this study, we refer to the Sharpe ratio model and reference (Khan et al., 2021) to transform the original constrained optimization model into an unconstrained single-objective optimization problem. The transformed objective function is:
(12) $F(E)=\max$ $\;\beta_{1}\left[\frac{\sum_{i=1}^{M}E_{i}\alpha_{i}-R}{\sqrt{\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}}}\right]-\beta_{2}\left(\sum_{i=1}^{M}E_{i}-1\right)^{2}$
@@ 2.4. Our unconstrained model
where $\left(\sum_{i=1}^{M}E_{i}-1\right)^{2}$ represents the equality constraint specified in Equation (10), ensuring the sum of the investment weights equals 1. While $(E_{i}-1\geq 0)(E_{i}-1)$ and $(-E_{i}\geq 0)(-E_{i})$ correspond to the inequality constraints outlined in Equation (11), which ensure the weights remain within valid bounds. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ denote the penalty weights associated with these constraints.
3. Methodology Overview: original DE and Proposed Improvements
In this section, we will first provide a detailed introduction to the original DE. Subsequently, we will conduct an in-depth study of the proposed QHDE, including a good point set-chaos reverse learning initialization mechanism, a quantum evolution strategy based on population information exchange, and a dynamic elite pool and Cauchy-Gaussian mixed disturbance strategy.
3.1. The original DE
@@
(14) $u_{i,j}(t+1)=\begin{cases}v_{i,j}(t+1)&\text{if }\text{rand}_{j}(0,1)\leq CR\text{ or }j=j_{\text{rand}}\\ x_{i,j}(t)&\text{otherwise}\end{cases}$
@@
(15) $x_{i}(t+1)=\begin{cases}u_{i}(t+1)&\text{if }f(u_{i}(t+1))\leq f(x_{i}(t))\\ x_{i}(t)&\text{otherwise}\end{cases}$
@@
3.2. Our proposed QHDE
The portfolio optimization involves multiple variables and complex constraints, requiring algorithms to efficiently explore large-scale search spaces. Therefore, the goal of the improvements is to enhance the algorithm’s performance in high-dimensional complex problems, especially in terms of stability and accuracy when dealing with multiple local optima and nonlinear constraints. To achieve this, quantum behavior, adaptive mechanisms and more efficient perturbation strategies are incorporated for strengthening the algorithm’s global search capability. These enhancements significantly boost the efficiency of the DE algorithm while mitigating the risk of being trapped in local optima. As a result, the algorithm becomes more robust and better equipped to handle the challenges of complex portfolio optimization problems.
3.2.1. a good point set-chaos reverse learning initialization mechanism
In metaheuristic algorithms, the quality of the initial solution significantly influences the search outcome. Since the random numbers generated by the system are not entirely random, they often exhibit disadvantages such as uneven distribution or tendency to cluster. Hence, to ensure that the values are as uniformly distributed as possible in the search space and to improve the ergodicity of the initial solution, we utilize the good points set to initialize the population. Let $G_{s}$ represent the unit hypercube in the $s$-dimensional space. If $r\in G_{s}$, then the set of points is given as:
(16) $GP_{n}(k)=\left(\left\{r_{1}^{(n)}\times k\right\},\ldots,\left\{r_{s}^{(n)}\times k\right\}\right),\quad 1\leq k\leq n$
@@ 3.2. Our proposed QHDE
Reverse learning is a perturbation mechanism, whose fundamental principle lies in obtaining the reverse solution of a known solution in the target space through reverse learning, and then choosing a superior solution from them as the current solution. Chaotic systems are a kind of nonlinear and highly sensitive deterministic systems. Although their behaviors do not exhibit periodicity or tend toward a stable state, they always maintain certain boundaries. The inherent regularity and randomness of chaotic systems, along with their ergodicity and unpredictability, make them a reliable source of randomness, offering more effective search strategies for heuristic optimization algorithms compared to traditional random sequences. In this paper, we adopt the Logistic chaotic mapping to initialize the population X generated by the optimal point set:
(18) $G^{t+1}=\mu\times G^{t}(1-G^{t})$
@@
Figure 1. Flow chart of QHDE
Flow chart of QHDE
3.2.2. quantum evolution strategy based on population information exchange
Traditional DE often encounters issues like slow convergence and a propensity to become stuck in local optima, particularly when addressing complex optimization problems like portfolio optimization. To address these challenges, this paper proposes integrating quantum theory into the DE framework. The quantum search mechanism simulates the process in which particles in a potential field move towards the lowest potential energy point in quantum mechanics. Specifically, the search space is considered as a potential field in quantum mechanics, the global optimum solution corresponding to the lowest point of the potential energy in the field, and the optimization process is the movement of particles towards this lowest potential energy point in this potential field. In quantum mechanics, the dynamic behavior of particles is described by the Schrödinger equation:
(20) $j\hbar\frac{\partial}{\partial\tau}\Psi(q,\tau)=\left(-\frac{\hbar^{2}}{2m}\nabla^{2}+V(q)\right)\Psi(q,\tau)$
where $\hbar$ represents the Planck constant, wave function $\Psi(q,\tau)$ represents the probability density of a particle being at position $q$ at time $\tau$, $m$ represents the mass of the particle, $\nabla^{2}$ represents the Laplacian operator, which represents the spatial distribution of the wave function, and $V(q)$ denotes the potential energy distribution function. The potential energy distribution of the Delta potential well is:
(21) $V(q)=-\gamma\delta(q)$
where $\gamma$ represents the depth of the potential well, and $\delta(q)$ denotes the Dirac delta function. The particle’s wave function is:
(22) $\Psi(q)=\frac{1}{\sqrt{L}}e^{-|q|/L}$
where $L$ represents the characteristic length parameter of the potential well, which controls the decay rate of the wave function. The probability density function is given by:
(23) $H(y)=|\Psi(y)|^{2}=\frac{1}{\sqrt{L}}e^{-2|q|/L}$
We use the Monte Carlo method to measure the position by collapsing the quantum state to a classical state:
(24) $y=\pm\frac{L}{2}\ln\left(\frac{1}{r_{4}}\right)$
where $r_{4}$ represents a random variable in $[0,1]$. Based on the above quantum theory, we propose a Quantum Evolution Strategy based on Population Information Exchange. The core idea of this strategy is to enhance the global search capability of the optimization algorithm by introducing a population information exchange mechanism and combining the characteristics of quantum evolution. First, the position of the population’s centroid is defined as:
(25) $x_{\text{mean}}=\frac{1}{N}\sum_{i=1}^{N}{x_{i}}$
The adaptive factor is defined as:
(26) $\eta=(w_{2}-w_{1})\times\frac{T_{\max}-t}{T_{\max}}+w_{1}$
where $w_{1}$ and $w_{2}$ denote control factors that determine the adjustment range of the adaptive factor, $T_{\max}$ represents the maximum number of iterations, and $t$ denotes the current iteration count. The parameter expressions are:
(27) $d=x_{i}+(c_{1}\times r_{5}\times(x_{i}-x_{\text{worse}}))+(c_{2}\times(1-r_{5})\times(x_{\text{best}}-x_{i}))$
(28) $b=\frac{\eta}{2}\times|x_{i}-x_{\text{best}}|$
(29) $z=\log\left(\frac{1}{r_{6}}\right)$
[…]
(30) $x_{new}=\begin{cases}d+b\times z,&\text{if }\text{rand}<0.5\\ d-b\times z,&\text{otherwise}\end{cases}$

### v3

# Equal contribution. Corresponding authors

@@ Abstract
High-dimensional portfolio optimization faces significant computational challenges under complex constraints, with traditional optimization methods struggling to balance convergence speed and global exploration capability. To address this, firstly, we introduce an enhanced Sharpe ratio-based model that incorporates all constraints into the objective function using adaptive penalty terms, transforming the original constrained problem into an unconstrained single-objective formulation. This approach preserves financial interpretability while simplifying algorithmic implementation. To efficiently solve the resulting high-dimensional optimization problem, we develop a Quantum Hybrid Differential Evolution (QHDE) algorithm, which introduces a dynamic quantum tunneling mechanism that enables individuals to probabilistically escape local optima, dramatically enhancing global exploration and solution flexibility. To further improve performance, a good point set–chaos reverse learning strategy generates a well-dispersed initial population, providing a robust and diverse starting point. Meanwhile, a dynamic elite pool combined with Cauchy-Gaussian hybrid perturbations maintains population diversity and mitigates premature convergence, ensuring stable and high-quality solutions. Experimental validation on CEC benchmarks and real-world portfolios involving 20 to 80 assets demonstrates that QHDE’s performance improves by up to 96.6%. It attains faster convergence, higher solution precision, and greater robustness than seven state-of-the-art counterparts, thereby confirming its suitability for complex, high-dimensional portfolio optimization.
1 Introduction
@@
In finance, a portfolio represents a collection of stocks or assets managed by individuals or institutions. Portfolio optimization entails distributing a fixed budget among the assets to minimize risk while maximizing expected returns. The Modern Portfolio Theory (MPT) employs the mean-variance model to describe the relationship between risk and return [9]. The core principle of this theory quantifies risk using variance, where lower-risk portfolios are typically characterized by smaller variances. However, MPT overlooks practical market constraints and the influence of background risks, such as base requirements, upper and lower bounds, and challenges inherent in managing large-scale asset portfolios [6, 10, 15]. Therefore, real-world portfolio optimization must incorporate these constraints and uncertainties to account for actual market conditions.
@@
Differential Evolution (DE) [3], as a classic heuristic optimization algorithm, is widely applied in various engineering optimization and scientific research tasks due to its simple implementation, few parameters, and strong global search capability [13, 22, 21]. Given the hybrid decision-making nature of discrete asset selection and continuous weight allocation in portfolio optimization, the real-valued encoding mechanism of DE facilitates a precise characterization of the continuous search space of asset weights, and combined with specialized mapping strategies to effectively handle complex constraints and multi-objective optimization, thereby adapting to both nonlinear and multi-modal requirements in financial decision-making. However, when confronted with high-dimensional complex portfolio problems, DE faces notable limitations. First, it is prone to becoming trapped in local optima, hindering the search for the global optimum. Second, as the dimensionality increases, the search efficiency declines and convergence slows. These limitations restrict the application effectiveness of DE in the field of portfolio optimization.
To overcome the aforementioned shortcomings, researchers have attempted to enhance the population diversity of the algorithm by introducing quantum-inspired mechanisms such as wave functions, probability amplitudes, and qubit representations [19, 17]. For instance, the evolution of probability density driven by the Schrödinger equation [14], operator reconstruction based on qubit representations [2], and the probability amplitude directional collapse induced by the quantum rotation gate [20] have been proven to significantly improve the global exploration efficiency of the algorithm. However, existing quantum-inspired operators often lack dynamic state awareness, leading to poor search precision in later iterations. Additionally, they struggle to balance energy barrier penetration with local exploitation when facing objective functions with complex rotation and shift characteristics.
@@
In response to the above challenges, we propose a variant of the DE based quantum theory, referred to as the Quantum Hybrid Differential Evolution Algorithm (QHDE). QHDE incorporates multiple innovative improvements relative to conventional DE. Inspired by the quantum tunneling behavior, QHDE constructed a dynamic quantum potential well model based on fitness differences. This enables individuals to bypass the local traps that were traditionally difficult for classical DE to overcome. Furthermore, the incorporation of chaotic reverse learning extends the coverage of initial solutions, resulting in a more advantageous search starting point. Meanwhile, the dynamic elite pool and Cauchy-Gaussian mixed disturbance strategy effectively prevent premature convergence and increase the diversity of the solutions. The combination of these improvement strategies enables QHDE to surpass traditional DE and other metaheuristic algorithms in portfolio optimization, allowing for faster identification of the optimal solution that balances risk and return. The main contributions of this paper are as follows:
(1) We develop a quantum tunneling strategy that transcends classical search limitations. By enabling non-classical probabilistic transitions, this strategy allows individuals to bypass local optima through a ’tunneling’ effect rather than stochastic jumps, fundamentally enhancing global exploration efficiency.
@@
(2) We introduce an initialization mechanism combining the good point set method with chaotic reverse learning. This enhances the quality of initial solutions, offering a more efficient starting point for subsequent optimization.
@@
(3) We propose a dynamic elite pool strategy and applies Cauchy-Gaussian random perturbation. This strategy mitigates premature convergence to local optima, enhances solution diversity, and strengthens global search capability.
@@
(4) We apply QHDE to portfolio optimization, where it outperforms seven state-of-the-art algorithms, exceeding the runner-up by up to 96.6%. This approach significantly accelerates convergence while enhancing accuracy, proving to be an efficient and robust solution for high-dimensional problems.
2 Math Equations
@@
This section introduces modeling methods of the portfolio selection problem, laying the foundation for the subsequent introduction of optimization techniques.
2.1 Classical Models
Mean-Variance Model: The objective function minimizes total portfolio risk, treating the expected return as a constraint. This ensures the portfolio meets a specific return threshold while maintaining the lowest possible risk level:
$\min\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}$ (1)
$\text{s.t.}\sum_{i=1}^{M}E_{i}\alpha_{i}=R$ (2)
@@ 2.1 Classical Models
where, $M$ denotes the total number of stocks, $E_{i}$ and $E_{j}$ represents the investment weight of the $i$th and $j$th stock respectively, $Q_{ij}$ signifies the covariance between the $i$th and $j$th stocks, and $\alpha_{i}$ denotes the expected return of the $i$th stock. Equation (2) ensures that the total return of the investment portfolio reaches the predetermined expected return $R$.
@@
Efficient Frontier: Unlike the mean-variance model, which treats expected returns as a constraint, the efficient frontier model integrates them directly into the objective function. It optimizes the risk-return trade-off by applying a penalty term to the return, and the expression is:
$\min\left[\omega\left(\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}\right)-(1-\omega)\left(\sum_{i=1}^{M}E_{i}\alpha_{i}\right)\right]$ (3)
where $\omega\in[0,1]$.
Sharpe Ratio Model: The Sharpe Ratio Model shifts focus from individual parameters to their ratio, measuring excess return per unit of risk. This single metric comprehensively balances portfolio performance and risk, expressed as:
$Sharpe-ratio=\frac{R_{p}-R}{std(p)}$ (4)
@@
$\max\frac{\sum_{i=1}^{M}E_{i}\alpha_{i}-R}{\sqrt{\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}}}$ (5)
In all the above models, the investment weights $E_{i}$ are subject to full allocation ($\sum E_{i}=1$) and non-negativity ($0\leq E_{i}\leq 1$) constraints. This ensures that the funds are fully allocated and no short-selling is allowed.
2.2 Our unconstrained model
In this study, we refer to the Sharpe ratio model and reference [6] to transform the original constrained optimization model into an unconstrained single-objective optimization problem. The transformed objective function is:
$F(E)=\max$ $\;\beta_{1}\left[\frac{\sum_{i=1}^{M}E_{i}\alpha_{i}-R}{\sqrt{\sum_{i=1}^{M}\sum_{j=1}^{M}E_{i}E_{j}Q_{ij}}}\right]-\beta_{2}\left(\sum_{i=1}^{M}E_{i}-1\right)^{2}$ (6)
@@ 2.2 Our unconstrained model
where $\left(\sum_{i=1}^{M}E_{i}-1\right)^{2}$ represents the equality constraint specified, ensuring the sum of the investment weights equals 1. While $(E_{i}-1\geq 0)(E_{i}-1)$ and $(-E_{i}\geq 0)(-E_{i})$ correspond to the inequality constraints outlined, which ensure the weights remain within valid bounds. $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ denote the penalty weights associated with these constraints.
3 Methodology Overview: original DE and Proposed Improvements
In this section, we first detail the original DE, then introduce the proposed QHDE, focusing on: a good point set-chaos reverse learning initialization, a potential-driven dynamic quantum tunneling strategy, and a dynamic elite pool with Cauchy-Gaussian mixed disturbance.
3.1 The original DE
@@
$u_{i,j}(t+1)=\begin{cases}v_{i,j}(t+1)&\text{if }\text{rand}_{j}(0,1)\leq CR\text{ or }j=j_{\text{rand}}\\ x_{i,j}(t)&\text{otherwise}\end{cases}$ (8)
@@
$x_{i}(t+1)=\begin{cases}u_{i}(t+1)&\text{if }f(u_{i}(t+1))\leq f(x_{i}(t))\\ x_{i}(t)&\text{otherwise}\end{cases}$ (9)
@@
3.2 Our proposed QHDE
Portfolio optimization in high-dimensional spaces presents significant challenges due to nonlinear constraints and numerous local optima. To address these, we enhance DE by integrating quantum tunneling, adaptive mechanisms, and efficient perturbation strategies. These improvements strengthen global search capabilities and mitigate premature convergence. Consequently, the proposed algorithm demonstrates superior robustness and precision, effectively navigating complex search spaces to achieve stable, high-quality solutions for large-scale portfolio problems.
3.2.1 A good point set-chaos reverse learning initialization mechanism
In metaheuristic algorithms, the quality of the initial solution significantly influences the search outcome. Standard pseudo-random generators often produce uneven distributions or clusters. Hence, to ensure maximum uniformity and ergodicity across the search space, we utilize the good points set to initialize the population. In the $s$-dimensional unit hypercube $G_{s}$, the set of points for $r\in G_{s}$ is defined as:
$GP_{n}(k)=\left(\left\{r_{1}^{(n)}\times k\right\},\ldots,\left\{r_{s}^{(n)}\times k\right\}\right),\quad 1\leq k\leq n$ (10)
@@ 3.2.1 A good point set-chaos reverse learning initialization mechanism
Reverse learning is a perturbation mechanism that compares a known solution with its reverse counterpart in the target space, selecting the superior one. Chaotic systems, known for their sensitivity, ergodicity, and unpredictability, provide a more robust source of randomness than traditional sequences. In this paper, we adopt the Logistic chaotic mapping to initialize the population X generated by the optimal point set:
$G^{t+1}=\mu\times G^{t}(1-G^{t})$ (12)
@@
3.2.2 Potential-driven dynamic quantum tunneling strategy
Traditional DE often suffers from slow convergence and local optima stagnation, especially in complex tasks like portfolio optimization. To address these challenges, this paper proposes integrating quantum tunneling theory into the DE framework. This mechanism models a individuals trapped in local optima as a particle confined within a time-variant potential well, defined by center $C$ and characteristic length $L$. As iterations $t$ progress, the potential well expands, enabling a non-classical probabilistic transition. Guided by the tunneling probability $P_{t}(X_{i})$, solutions can "tunnel" through energy barriers (fitness gaps) to relocate beyond local basins. This probabilistic mechanism significantly boosts global exploration and the algorithm’s ability to reach the global optimum. The strategy is structured into three primary components:
1. Potential Energy Modeling: The fitness difference of an individual $x_{i}$ is mapped to a quantum potential barrier $V(x_{i})$, defined as follows:
$V(x_{i})=\frac{|fit_{i}-f_{\text{best}}|}{\max(f)-\min(f)}\cdot V_{0}$ (14)
where $f_{\text{best}}$ denotes the current global best fitness, and $V_{0}$ denotes the baseline barrier height. This model ensures that individuals closer to the optimum face lower potential barriers, while those further away encounter higher resistance.
2. Tunneling Probability Calculation: Based on the Wentzel-Kramers-Brillouin (WKB) approximation, the equivalent tunneling probability $P_{t}(x_{i})$ is formulated to determine the likelihood of an individual penetrating the "energy barrier":
$P_{t}(x_{i})=\exp\left(-\frac{2\sqrt{2m}}{\hbar}\int_{a}^{b}\sqrt{V(x)-E_{k}}\,dx\right)$ (15)
where $E_{k}=\frac{1}{N}\sum_{i=1}^{N}f(x_{i})$ represents the average kinetic energy of the population, $\gamma\in[0.05,0.2]$ denotes the tunneling adjustment coefficient. When the potential energy $V(x_{i})$ is lower than $E_{k}$, $P_{t}$ is set to 1, facilitating free movement. Conversely, if $V(x_{i})>E_{k}$, the probability decreases exponentially, allowing for controlled, non-classical probabilistic jumps.
3. Intelligent Position Update: If the tunneling condition ($rand<P_{t}(x_{i})$) is triggered, the individual’s position is updated via a coupled perturbation mechanism:
$x_{\text{new}}=x_{i}+\beta\cdot T(t)\cdot\Delta(t)$ (16)
In this update rule, $\beta\sim\mathcal{N}(0,1)$ follows the standard normal distribution, providing a random search direction. While the time-varying coefficient $T(t)$ and the dynamic step size $\Delta(t)$ collaboratively balance global exploration in early stages with local fine-tuning in later iterations.